OpenStudy (anonymous):

suppose x varies directly as y and inversely as z, and x=96 when y=8 and z=12. find x when y=4 and z=6

OpenStudy (anonymous):

x=(ay)/z 96=(a*8)/12 a=144 when y=4,z=6 then x=(144*4)/6=96